using the domain values {-1,0,4}, find the range values for relation y=2x-10
Answer by Edwin McCravy(20059)

The domain is the set of x-values.  They range is the set of y-values
[You can remember which is which because "d" for "domain" comes before "r" 
for "range" in the alphabet, and "x" also comes before "y" in the alphabet.] 

using the domain x-values {-1,0,4}, find the range y-values for relation
y=2x-10
To find the range y-value for the domain x-value -1, 
substitute -1 for x in the equation, and solve for y:

y = 2x - 10
y = 2(-1) - 10
y = -2 - 10
y = -12

So the range y-value corresponding to the domain x-value -1 is -12.

To find the range y-value for the domain x-value 0, 
substitute 0 for x in the equation, and solve for y:

y = 2x - 10
y = 2(0) - 10
y = 0 - 10
y = -10

So the range y-value corresponding to the domain x-value 0 is -10.

To find the range y-value for the domain x-value 4, 
substitute 4 for x in the equation, and solve for y:

y = 2x - 10
y = 2(4) - 10
y = 8 - 10
y = -2

So the range y-value corresponding to the domain x-value 4 is -2.

So the set of range y-values corresponding to the domain x-values {-1,0,4},
for the relation y=2x-10 is {-12,-10,-2}.

Edwin